A 4.66 L sample of gas is warmed from 273 K to a final temperature of 398 K. Assuming no change in pressure, what is the final volume of the gas?
Answer #1

Given:

  • Initial volume (V1) = 4.66 L

  • Initial temperature (T1) = 273 K

  • Final temperature (T2) = 398 K

  • Pressure constant

Formula (Charles's Law):

V1T1=V2T2

Solve for final volume (V2):

V2=V1×T2T1=4.66L×398K273K

Calculation:

V2=4.66×1.45786.79L

Answer:
The final volume is 6.79 L.


answered by: anonymous